A man saw three girl in a field where they found a pile of apples and a boy. They decide to sleep in the field overnight and divide the pile ofapple in the morning. But during the night, one girl wakes up and gives one apple to the boy, and takes exactly 1/3 of the rest of the apples and went back to sleep. Then the second girl waked up and does the same thing . Later the third girl wakes up and does the same. In the morning they were fewer than 10 apples left. They each take 1/3. How many apples were there in the original pile.

Answer by venugopalramana(3286)

A man saw three girl in a field where they found a pile of apples and a boy. They decide to sleep in the field overnight and divide the pile ofapple in the morning. But during the night, one girl wakes up and gives one apple to the boy, and takes exactly 1/3 of the rest of the apples and went back to sleep. Then the second girl waked up and does the same thing . Later the third girl wakes up and does the same. In the morning they were fewer than 10 apples left. They each take 1/3. How many apples were there in the original pile.
LET THE APPLES NUMBER BE = X
I GIRL
ONE GIVEN TO BOY ..
REMAINING =X-1
TAKEN BY GIRL 1/3...=(X-1)/3
REMAINING = (X-1) - (X-1)/3 = 2(X-1)/3
------------------------------------
II GIRL
ONE GIVEN TO BOY ..
REMAINING =2(X-1)/3 -1 = (2X-5)/3
TAKEN BY GIRL 1/3...=(2X-5)/9
REMAINING = (2X-5)/3 - (2X-5)/9 = 2(2X-5)/9
-------------------------------------
III GIRL
ONE GIVEN TO BOY ..
REMAINING =2(2X-5)/9 - 1 = (4X-19)/9
TAKEN BY GIRL 1/3...=(4X-19)/27
REMAINING = (4X-19)/9 - (4X-19)/27 = 2(4X-19)/27
------------------------------------------------------
2(4X-19)/27 <10...THIS HAS TO BE A MULTIPLE OF 3 SO IT COULD BE 9 0R 6 0R 3
TAKE 9 FIRST
4X-19=270/2.GIVES FRACTIONS ..SO NOT POSSIBLE
TAKE 6
4X-19=81
4X=100
X=25.............OK
X<154/4=38.5
THIS HAS TO BE A MULTIPLE OF 3